OpenStudy (anonymous):

Sam has 3 numbers \(\large a,b,c>0\). Sam has stated that the numbers follow the condition: \(\large ab+bc+ac=abc\) Adi has been assigned the task of finding the smallest possible value of \(\large a+b+c\). Can you help him by finding this value?

OpenStudy (anonymous):

Hint: apply AM-HM

OpenStudy (anonymous):

\(\large \color{blue}{Solution} \) Given numbers are positive. We can always think of applying AM GM HM inequality. Note that we have \(\large ab+bc+ac=abc\) Now, what we can think here is dividing both sides by abc nd we get \(\large \frac{1}{a}+\frac{1}{b}+\frac{1}{c}=1\) Success. Now we can directly apply AM HM. Nd we obtain \(\large (a+b+c)(\frac{1}{a}+\frac{1}{b}+\frac{1}{c} )\ge9\) Simply giving minimum of \(\large a+b+c\) equal to 9